OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 2x + 5 A. (15, 5) and (2, 9) B. (5, 10) and (2, 9) C. (5, 15) and (9, 2) D. (5, 15) and (2, 9)

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

Hey!

OpenStudy (anonymous):

We need to plug in the points into the equation and see if they work!

OpenStudy (anonymous):

I'll do one point as an example for you!! (15,5) That is in the form of (x,y) x = 15 y = 5 Plug it in to the equation! y = 2x + 5 5 = 2(15) + 5 5 = 30 + 5 5 = 35 That is not true, so (15,5) is not a solution to the equation! That means we can eliminate any answers that has (15,5)

OpenStudy (sushi121212):

Just plug in the numbers for x and y: (5, 15) and (2, 9) The 5 and the 2 is a x, and the 15 and the 9 is a y. y = 2x + 5 15=2(5)+5 15=10+5 15=15 If it equals each other, then that means that it is the answer. y = 2x + 5 9=2(2)+5 9=4+5 9=9 Your answer is (5, 15) and (2, 9)!

OpenStudy (anonymous):

@Sushi121212 Doing all the work and giving the asker doesn't help them learn. :c Please read the Code of Conduct which states the rules that OpenStudy users must follow. http://openstudy.com/code-of-conduct

OpenStudy (sushi121212):

I'm sorry, I just wanted to help.